ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 87, 1994

Aaronson, J. and M. Keane Isomorphism of random walks  37

Aaronson, J., T. Liggett and P. Picco Equivalence of renewal sequences and isomorphism of random walks

65

Abramenko, P. Finiteness properties of Chevalley groups over 𝔽_{q}[t] 203
Amberg, B. and A. Fransman Products of groups and group classes  9
Amitsur, S. A. and L. H. Rowen Elements of reduced trace 0  161
Bourgain, J. Hausdorff dimension and distance sets  193
Duval, A. M. A combinatorial decomposition of simplicial complexes  77
Farmer, J. D. Polynomial reflexivity in Banach spaces  257

Fransman, A. See B. Amberg

Gabber, O. Affine analog of the proper base change theorem  325

Gaboriau, D., G. Levitt and F. Paulin Pseudogroups of isometries of ℝ and Rips’ theorem on free actions on ℝ-trees

403

Gluskin, E., M. Meyer and A. Pajor Zeros of analytic functions and norms of inverse matrices

225

Gorelik, E. The uniform nonequivalence of L_{p} and l_{p}  1

Gowers, W. T. A finite-dimensional normed space with two nonequivalent symmetric bases

143

Grekas, S. Structural properties of compact groups with measuretheoretic applications

89

Grønbæk, N., B. E. Johnson and G. A. Willis Amenability of Banach algebras of compact operators

289

Huberich, M. Non-regular ultrafilters  275

Johnson, B. E. See N. Grønbæk

Johnson, W. B. and G. Schechtman Computing p-summing norms with few vectors

19

Keane, M. See J. Aaronson

Kudla, S. S. Splitting metaplectic covers of dual reductive pairs  361

Kwiatkowski, J., M. Lemańczyk and D. Rudolph A class of real cocycles having an analytic coboundary modification

337

Lemańczyk, M. See J. Kwiatkowski

Leung, D. H. Some isomorphically polyhedral Orlicz sequence spaces 117

Levitt, G. See D. Gaboriau Liggett, T. See A. Aaronson Lind, D. and K. Schmidt Bernoullicity of solenoidal automorphisms

and global fields

33

Meyer, M. See E. Gluskin

Notbohm, D. Kernels of maps between classifying spaces  243

Pajor, A. See E. Gluskin Paulin, F. See D. Gaboriau Picco, P. See J. Aaronson Rowen, L. H. See S. A. Amitsur Rudolph, D. See J. Kwiatkowski Schechtman, G. See W. B. Johnson Schmidt, K. See D. Lind

Shalev, A. A new characterization of Frobenius complements  153

Shumyatsky, P. On periodic solvable groups having automorphisms with nilpotent fixed point groups

111

Talagrand, M. Orlicz property and cotype in symmetric sequence spaces

181

Willis, G. A. See N. Grønbæk Yang, S. Estimates for coefficients of univalent functions from integral

means and Grunsky inequalities

129

Ziegler, G. M. Shellability of chessboard complexes  97